one airplane pilot flew 270 miles per hour to her destination. On her return trip, she flew at 360 miles per hour. Her total flying time was 7 hours. How far was it from one airport to another?
Answer by mananth(16946)

time = distance/speed
time forward + time return = total time =7 hours
Let distance be x miles
x/270 + x/360 = 7 hours
360X+270x/(270*360)=7
630x=7*270*360
x=7*270*360/630
x= 1080 miles
...
m.ananth@hotmail.ca